OpenStudy (anonymous):

i need the solution set for sin(x)=1/2, where x is positive

OpenStudy (anonymous):

sin of 30deg or sin 150deg

OpenStudy (nincompoop):



OpenStudy (anonymous):

the full question: \[\exists! x x \in[0,\theta] and \sin(x)=1/2\]

OpenStudy (nincompoop):

take the inverse of both sides

OpenStudy (nincompoop):

take the inverse sine of both sides

OpenStudy (anonymous):

Sin(x) is 1/2 when x is a coterminal of pi/6 or 5pi/6, so the set S is pi/6 + 2pi(n) unioned with 5pi/6 +2pi(n), where n is a non-negative integer.